A researcher investigated the number of viral infections people contracted as a function of the amount of stress they experienced during a 6-month period. She obtained the following data:

Report the means which significantly differ from one another according to the Tukey post-hoc results?

Negligible Stress Minimal Stress Moderate Stress Severe Stress
2 4 6 5
1 3 5 7
4 2 7 8
1 3 5 4
Answer #1
Anova: Single Factor
SUMMARY
Groups Count Sum Average Variance std dev
Negligible Stress 4 8 2.000 2.0 1.4
Minimal Stress 4 12 3.000 0.7 0.8
Moderate Stress 4 23 5.750 0.9 1.0
4 24 6.00 3.33
ANOVA
Source of Variation SS df MS F P-value F crit
Between Groups 47.688 3 15.90 9.19 0.0020 3.49
Within Groups 20.750 12 1.73
Total 68.438 15
Level of significance 0.05
no. of treatments,k 4
DF error =N-k= 12
MSE = 1.7292
q-statistic value(α,k,N-k) 4.1985

critical value = q*√(MSE/n) = 2.76

if absolute difference of means > critical value,means are significnantly different ,otherwise not

population mean difference critical value result
µ1-µ2 1.00 2.76 means are not different
µ1-µ3 3.75 2.76 means are different
µ1-µ4 4.00 2.76 means are different
µ2-µ3 2.75 2.76 means are not different
µ2-µ4 3.00 2.76 means are different
µ3-µ4 0.25 2.76 means are not different